# Maximum-likelihood estimation of the Matérn covariance structure of isotropic spatial random fields on finite, sampled grids

### v1

@@ 2.2.3 Simulation
Eq. (23) shows that the covariance of a finite set of gridded Fourier coefficients suffers both from blurring by the sampling kernel, as we have just illustrated and calculated explicitly, but also from correlation between the wavenumbers. To prepare for what is coming, we note, first, that the off-diagonal terms in $\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k^{\prime}})\rangle$ decay rapidly enough in most cases for us to ignore them as the basis for the simulations that we just showed, which took $\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k})\rangle\approx\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ as their point of departure, whether calculated on the interior domain of a doubled grid, using grid refinement, discrete convolution, and subsampling to approximate $\bar{\mathcal{S}}_{\boldsymbol{\theta}}$, or exactly, via eq. (30). Second, we will show empirically that we are able to ignore them when designing the debiased-Whittle likelihood [Guillaumin et al.(2017)Guillaumin, Sykulski, Olhede, Early, & Lilly, Sykulski et al.(2019)Sykulski, Olhede, Guillaumin, Lilly, & Early] to perform parameter estimates from sampled data, which is a central feature of this paper and its predecessors. Finally, we show that they will, however, play an important role in the calculation of the estimation variance of the maximum-likelihood estimates, using the results obtained by [Guillaumin et al.(2022)Guillaumin, Sykulski, Olhede, & Simons], and we discuss various algorithms to conduct the relevant calculations. This last fact stands in apparent contradiction to the material discussed by [Simons & Olhede(2013)], their Sections A6 and A8, which, in retrospect, have proven to be overly optimistic, asymptotically.
As to data simulation, we are, alternatively, able to take wavenumber correlation into account also by switching to space-domain methods that use the spatial covariance, eq. (15), as their point of departure, via circulant embedding of the covariance matrix [Kroese & Botev(2015)].
@@ 3 W H I T H E R  W H I T T L E  ?
In this paper we develop a maximum-likelihood procedure that takes gridded input ‘topographies’ and estimates the three-element sets $\boldsymbol{\theta}$, see eq. (13), that contain the parameters of the isotropic Matérn spectral densities ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ or spatial covariances ${\mathcal{C}}_{\boldsymbol{\theta}}(r)$ by which we aim to sufficiently describe such planetary data patches. Before proceeding, we take a brief detour to illustrate, for the example of the variance, $\sigma^{2}$}, why we embark on this journey. Additional motivation and considerations are offered by, among others, [Vanmarcke(1983)] and [Stein(1999)].
@@
Eq. (35) applies quite generally, to stationary processes with spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ or covariance function ${\mathcal{C}}_{\boldsymbol{\theta}}(\mathbf{x})$. The error made in reducing the last term in (35) to the second term in eq. (36) should be interpreted as arising from the discretization of eq. (7). The appearance of the blurred spectral density, $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(0)$, is due to the finite-sample effects by which the spatial grid is relatively coarse, and too small to comprise the lags at which the structure is completely decorrelated. Only for, effectively, uncorrelated white noise, ${\mathcal{C}}_{\boldsymbol{\theta}}(\mathbf{x}-\mathbf{x^{\prime}})={\sigma^{2}}\delta(\mathbf{x},\mathbf{x^{\prime}})$, does eq. (35) reduce to the independent and identically distributed bias expression [Bendat & Piersol(2000)]
@@
Fig. 3 illustrates the arguments made so far in this section, for a variety of values of $\nu$ and $\rho$, as a function of grid size, and where the expectation of the estimate is formed by averaging over a great number of experiments. The naive variance estimator $s^{2}$ is biased, in a manner and for a reason that we understand intuitively, and are able to compute analytically. Rather than writing down expressions for the variance of the naive variance estimator $s^{2}$, we will illustrate its behavior on the basis of another suite of numerical experiments. Fig. 4 (left panels) reveals that the estimation variance is generally high (relatively speaking), and decaying too slowly (for our taste) with increasing grid size.
@@
In comparison, the maximum-likelihood estimator that we develop in the next section has properties that are far more favorable, as is readily, if proleptically, illustrated by the second suite of experiments shown in Fig. 4 (right panels, marked ‘MLE’). Saving the details of its construction for the next section, inspecting the figure reveals that, as soon as the data patch size exceeds the decorrelation length of the Matérn process, the estimation variance of the maximum-likelihood variance estimator is acceptably low. Moreover, the estimation variance continues to decay at a pleasing rate, suggestive of its asymptotic unbiasedness.
@@ 4 M A X I M U M - L I K E L I H O O D  T H E O R Y
The material in this section is chiefly inspired by sections 4.3–4.8 and Appendix A6 of [Simons & Olhede(2013)]. [Cox & Hinkley(1974)] remains an excellent reference, though more modern texts such as [Pawitan(2001)] and [Severini(2001)] are recommended. Our main device is the frequency-domain [Whittle(1953), Whittle(1954)] likelihood, modified to acknowledge edge effects by blurring the spectral density function.
@@ 4.1 Finite large-sample theory
[Simons & Olhede(2013)] introduced ${\mathcal{L}}(\boldsymbol{\theta})$, the likelihood for the Matérn parameters, which, as is acceptable only for large sample sizes, ignores the blurring in the isotropic spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ as well as the correlation induced between wavenumbers, in the form
@@
Its first derivatives with respect to each of the parameters $\theta$ are the elements of the unblurred score vector ${\mbox{\boldmath$\gamma$}}(\boldsymbol{\theta})$, given by
$\gamma_{\theta}(\boldsymbol{\theta})=-\frac{1}{MN}\sum_{\mathbf{k}}m_{\theta}(k)\left[1-{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}\right],$ (40)
with, easily obtained via differentiation of eq. (12), and listed as eqs (73) in the Appendix, the required expressions
@@
Its second derivatives are the elements of the Hessian matrix $\mathbf{F}(\boldsymbol{\theta})$,
$F_{\theta\theta^{\prime}}(\boldsymbol{\theta})=-\frac{1}{MN}\sum_{\mathbf{k}}\Bigg{[}\frac{\partial m_{\theta^{\prime}}\hskip-1.00006pt(k)}{\partial\theta}+\left\{m_{\theta}(k)m_{\theta^{\prime}}\hskip-1.00006pt(k)-\frac{\partial m_{\theta^{\prime}}\hskip-1.00006pt(k)}{\partial\theta}\right\}\left\{{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}\right\}\Bigg{]},$ (42)
@@
${\mathcal{F}}_{\theta\theta^{\prime}}(\boldsymbol{\theta})=\frac{1}{MN}\sum_{\mathbf{k}}m_{\theta}(k)\hskip 1.00006ptm_{\theta^{\prime}}\hskip-1.00006pt(k).$ (43)
@@
$\hat{\theta}-z_{\beta/2}\frac{\mathcal{J}_{\theta\theta}^{1/2}({\mbox{\boldmath$\hat{\theta}$}})}{\sqrt{MN}}\leq\theta_{0}\leq\hat{\theta}+z_{\beta/2}\frac{\mathcal{J}_{\theta\theta}^{1/2}({\mbox{\boldmath$\hat{\theta}$}})}{\sqrt{MN}}.$ (45)
The relations in this section are theoretical quantities derived by [Simons & Olhede(2013)] that, strictly speaking, apply only to the ‘population’ case, on domains of infinite extent. As mentioned before, however, in the ‘sample’ case of discretized, windowed likelihood analysis, we must replace the Matérn spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ with a suitably blurred version, $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(k)$, to acknowledge the effects of finite sampling. In what follows, we will explore the implications for the uncertainty estimates of the parameters.
@@ 4.2 Finite sampled data: heuristics
For sampled data, [Simons & Olhede(2013)] defined the likelihood of observing the data ${\mathcal{H}}(\mathbf{x})$ under the spectral model (12)–(13) parameterized by $\boldsymbol{\theta}$. This ‘blurred likelihood’ $\bar{{\mathcal{L}}}(\boldsymbol{\theta})$ instead of ${\mathcal{L}}(\boldsymbol{\theta})$, is given in terms of the Fourier coefficients of the gridded and windowed data, $H(\mathbf{k})$ in eq. (25), and of the blurred isotropic spectral density, $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ of eq. (30), summed over all wavenumbers in the grid (21), by
@@
Compare eq. (46) with eq. (39): the only difference is the acknowledgment of the spectral blurring effect of sampled data. Eq. (46) is the quantity that we maximize, under positivity constraints, for the parameter vector $\boldsymbol{\theta}$, thereby defining the quasi-maximum-likelihood estimate $\hat{\theta}$ to be the maximizer of the ‘score’ vector ${\mbox{\boldmath$\bar{\gamma}$}}(\boldsymbol{\theta})$ of numerical first derivatives of the blurred likelihood, which solves
${\mbox{\boldmath$\bar{\gamma}$}}({\mbox{\boldmath$\hat{\theta}$}})=\mathbf{0}.$ (47)
@@
Satisfying eq. (47) to find $\hat{\theta}$, for example by an iterative function minimization routine, requires repeated evaluation of the spectral density (12) on the grid (21), with the blurring implemented convolutionally (on a refined and subsequently reinterpolated grid) or else exactly, as discussed in Section 2.2.2. Without entering into more details for now, Fig. 5 shows the results of a suite of experiments conducted to illustrate the performance of our numerical method that recovers each of the three Matérn parameters $(\sigma^{2},\nu,\rho)$, as a function of grid size (measured in terms of the correlation length $\pi\rho$). As the right-hand side panels of Fig. 4 showed for the variance estimate, the maximum-likelihood estimates are very well-behaved, from about the point where the grid size reaches a linear dimension of about $\pi\rho$. The procedure almost surely yields low-variance and practically unbiased estimates from a grid size of about $2\pi\rho$ onward, as Fig. 5 shows empirically.
[…]
@@
The mathematical form and geometry of the blurred likelihood function are what, fundamentally, controls the observed behaviors. [Simons & Olhede(2013)] only considered convolutional approaches to blurring, and were limited in their ability to acknowledge the spectral interaction induced by the applied data windows on the parameter covariance estimates. The first and second derivatives of the blurred likelihood are not simply obtained by substituting $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(k)$ for ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ in eqs (40) and (42), since the factors $m_{\theta}(k)$ that appear in the expressions for the score and the Hessian are analytical derivatives that involve the unblurred spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$. Replacing ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ by $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(k)$ in eq. (40) yields reasonable approximations for the likelihood derivatives, which compare relatively favorably to their numerical counterparts—especially for large sample sizes. Most numerical optimization routines will be able to maximize eq. (46), solving eq. (47) without being given explicit expressions for the score and the Hessian. However, in order to derive accurate confidence intervals on our estimated parameters, we do need access to the expected value of the second-order derivatives of the actual likelihood that is being maximized.
@@
Numerical experiments and theoretical considerations along the lines suggested in their Appendix A8 tempted [Simons & Olhede(2013)] into concluding that eq. (45) could be used to construct confidence intervals for the solutions of eq. (47) in our present case of univariate two-dimensional analysis. Under the viewpoint (espoused in their eqs A84–A87), the blurred spectrum is an additive correction term (small for smoothly varying spectra) away from the original. In this framework, neglecting to blur the Fisher matrix (not to mention neglecting wavenumber correlations) was believed to have an altogether negligible effect on the estimation variances based on its inverse, even if blurring the likelihood is absolutely essential to arrive at the estimate in the first place. However, the ability of the unblurred Fisher matrix to help predict the variance of the parameters derived via maximization of the blurred likelihoods turns out to be poor, especially as concerns the variance and correlation parameters $\sigma^{2}$} and $\rho$. The unblurred expression, eq. (43), of the Fisher matrix provides an asymptotic but ultimately inadequate match to the average of the numerical Hessian for real-world sampling scenarios.
@@
Not accounting for wavenumber correlation proved to be another stumbling block. [Simons & Olhede(2013)] conceived of approximations to account for wavenumber correlation involving a multiplicative correction term (their eqs A56–A58). For very large sample sizes this correction term approaches unity. Contrary to the optimism they expressed, uncertainty estimates for the maximizers of eq. (46) that rely on eq. (45) are inadequate for all but the largest sample sizes. A heuristic way of determining the estimation variance for the recovered parameters when actual data are being investigated is to generate synthetics with features identical to those of the gridded data, from models with Matérn parameters given by previously obtained solutions, then estimating their parameters a number of times, and learning from their distribution what the likely uncertainty ranges for the parameters of the actual data patches might be, as in Fig. 5. However justified, little transferable knowledge is gained in the process, and the procedure is cumbersome and time-consuming.
@@
[Guillaumin et al.(2022)Guillaumin, Sykulski, Olhede, & Simons] showed the way forward by further developing the theory of likelihood analysis for finite sampled data, on which we rely to develop the practical methods offered in the next sections. They include the ability to calculate uncertainty estimates on the parameters from first principles. The next section provides a complete description of the entire workflow.

### v2

@@ 2.2.3 Simulation
Eq. (23) shows that the covariance of a finite set of gridded Fourier coefficients suffers both from blurring by the sampling kernel, as we have just illustrated and calculated explicitly, but also from correlation between the wavenumbers. To prepare for what is coming, we note, first, that if the off-diagonal terms in $\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k^{\prime}})\rangle$ decay rapidly enough, one could ignore them as the basis for simulations, taking $\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k})\rangle\approx\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ as a point of departure, whether calculated on the interior domain of a doubled grid, using grid refinement, discrete convolution, and subsampling to approximate $\bar{\mathcal{S}}_{\boldsymbol{\theta}}$ [[, as was done by]]Simons+2013, or exactly, via eq. (30). Here we do take wavenumber correlations into account for data simulation by switching to space-domain methods that use the spatial covariance, eq. (15), as their point of departure, via circulant embedding of the covariance matrix [41]. Second, we will show empirically that we are able to ignore them when designing the debiased-Whittle likelihood [26, 62] to perform parameter estimates from sampled data, which is a central feature of this paper and its predecessors. Finally, we show that they will, however, play an important role in the calculation of the estimation variance of the maximum-likelihood estimates, using the results obtained by [28], and we discuss various algorithms to conduct the relevant calculations. This last fact stands in apparent contradiction to the material discussed by [58], their Sections A6 and A8, which, in retrospect, have proven to be overly optimistic, asymptotically.
@@ 3 W H I T H E R  W H I T T L E  ?
In this paper we develop a maximum-likelihood procedure that takes gridded input ‘topographies’ and estimates the three-element sets $\boldsymbol{\theta}$, see eq. (13), that contain the parameters of the isotropic Matérn spectral densities ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ or spatial covariances ${\mathcal{C}}_{\boldsymbol{\theta}}(r)$ by which we aim to sufficiently describe such planetary data patches. Before proceeding, we take a brief detour to illustrate, for the example of the variance, $\sigma^{2}$}, why we embark on this journey. Additional motivation and considerations are offered by, among others, [63] and [61].
@@
Eq. (35) applies quite generally, to stationary processes with spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ or covariance function ${\mathcal{C}}_{\boldsymbol{\theta}}(\mathbf{x})$. The error made in reducing the last term in (35) to the second term in eq. (36) should be interpreted as arising from the discretization of eq. (7). The appearance of the blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(0)$ is due to the finite-sample effects by which the spatial grid is relatively coarse, and too small to comprise the lags at which the structure is completely decorrelated. Only for, effectively, uncorrelated white noise, ${\mathcal{C}}_{\boldsymbol{\theta}}(\mathbf{x}-\mathbf{x^{\prime}})={\sigma^{2}}\delta(\mathbf{x},\mathbf{x^{\prime}})$, does eq. (35) reduce to the independent and identically distributed bias expression [4]
@@
Fig. 3 illustrates the arguments made so far in this section, for a variety of values of $\nu$ and $\rho$, as a function of grid size, and where the expectation of the estimate is formed by averaging over a great number of experiments. The naive variance estimator $s^{2}$ is biased, in a manner and for a reason that we understand intuitively, and are able to compute analytically. Rather than writing down expressions for the variance of the naive variance estimator $s^{2}$, we will illustrate its behavior on the basis of another suite of numerical experiments.
@@
Fig. 4 (left panels) reveals that the estimation variance of $s^{2}$ is generally high (relatively speaking), and decaying too slowly (for our taste) with increasing grid size. In comparison, the quasi-maximum-likelihood estimator that we develop in the next section has properties that are far more favorable, as is readily, if proleptically, illustrated by the second suite of experiments shown in Fig. 4 (right panels, marked ‘MLE’). Saving the details of its construction for the next section, inspecting the figure reveals that, as soon as the data patch size exceeds the correlation length of the Matérn process, the estimation variance of the maximum-likelihood variance estimator is acceptably low. Moreover, the estimation variance continues to decay at a pleasing rate, suggestive of its asymptotic unbiasedness.
@@ 4 M A X I M U M - L I K E L I H O O D  T H E O R Y
The material in this section is chiefly inspired by sections 4.3–4.8 and Appendix A6 of [58]. [7] remains an excellent reference, though more modern texts such as [48] and [56] are recommended. Our main device is the frequency-domain [68, 69] likelihood, modified to acknowledge edge effects by blurring the spectral density function.
@@ 4.1 Finite large-sample theory
[58] introduced ${\mathcal{L}}(\boldsymbol{\theta})$, the likelihood for the Matérn parameters, which, as is at best acceptable only for large sample sizes, ignores the blurring in the isotropic spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ as well as the correlation induced between wavenumbers, in the form
@@
Its first derivatives, with respect to each of the parameters $\theta\in\boldsymbol{\theta}$, are the components of the unblurred score vector ${\mbox{$\gamma$}}(\boldsymbol{\theta})$, given by
$\gamma_{\theta}(\boldsymbol{\theta})=\frac{\partial{\mathcal{L}}(\boldsymbol{\theta})}{\partial\theta}=-\frac{1}{MN}\sum_{\mathbf{k}}m_{\theta}(k)\left[1-{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}\right],$ (40)
with, easily obtained via differentiation of eq. (12), and listed as eq. (73) in the Appendix, the required expressions
@@
Its second derivatives, with respect to two arbitrary parameters $\theta,\theta^{\prime}\in\boldsymbol{\theta}$, are the components of the Hessian matrix $\mathbf{F}(\boldsymbol{\theta})$,
$F_{\theta\theta^{\prime}}(\boldsymbol{\theta})=\frac{\partial\gamma_{\theta^{\prime}}(\boldsymbol{\theta})}{\partial\theta}=\frac{\partial^{2}{\mathcal{L}}(\boldsymbol{\theta})}{\partial\theta\partial\theta^{\prime}}=-\frac{1}{MN}\sum_{\mathbf{k}}\Bigg[\frac{\partial m_{\theta^{\prime}}\hskip-1.00006pt(k)}{\partial\theta}+\left\{m_{\theta}(k)m_{\theta^{\prime}}\hskip-1.00006pt(k)-\frac{\partial m_{\theta^{\prime}}\hskip-1.00006pt(k)}{\partial\theta}\right\}\left\{{\mathcal{S}}_{\boldsymbol{\theta}}^{-1}\hskip-1.00006pt(k)\hskip 1.00006pt|H(\mathbf{k})|^{2}\right\}\Bigg],$ (42)
@@
${\mathcal{F}}_{\theta\theta^{\prime}}(\boldsymbol{\theta})=-\left\langle\frac{\partial^{2}{\mathcal{L}}(\boldsymbol{\theta})}{\partial\theta\partial\theta^{\prime}}\right\rangle=\frac{1}{MN}\sum_{\mathbf{k}}m_{\theta}(k)\hskip 1.00006ptm_{\theta^{\prime}}\hskip-1.00006pt(k).$ (43)
@@
$\hat{\theta}-z_{\beta/2}\frac{\mathcal{J}_{\theta\theta}^{1/2}({\mbox{$\hat{\theta}$}})}{\sqrt{MN}}\leq\theta_{0}\leq\hat{\theta}+z_{\beta/2}\frac{\mathcal{J}_{\theta\theta}^{1/2}({\mbox{$\hat{\theta}$}})}{\sqrt{MN}}.$ (45)
The relations in this section are theoretical quantities derived by [58] that, strictly speaking, apply only to the ‘population’ case, on domains of infinite extent, and we note that caveats about consistency apply for $MN$ large in the fixed-domain asymptotic regime, as we alluded to in Sec. 1. While even under increasing-domain asymptotics, there may be few rigorous results on maximum likelihood estimates for the Matérn model when the smoothness parameter is treated as unknown, the (highly) significant correlation contribution introduced by [28] helps illuminate the estimability of, and the identifiability between, the three parameters, and the separation of different processes under various spatial sampling mechanisms, which will enable pursuing these issues further. In the ‘sample’ case of discretized, windowed likelihood analysis, we follow [58] and [28] in replacing the Matérn spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ in eq. (39) with a suitably blurred version, $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(k)$, to acknowledge the effects of finite sampling. In what follows, we will explore the implications of sampling and bounding on the uncertainty estimates of the parameters.
When using the Whittle likelihood as defined in eq. (39) to produce estimators for time series, the behavior of the spectral estimates is very well understood, see for example [15]. Exactly how the Whittle likelihood is formulated may vary—in terms of an integral or a sum, with the log covariance term included or not [37], and with and without tapering [12]. In two dimensions and higher, however, it is recognized that the boundary effects become dominant, which has led to alternative formulations [32, 14, 28]. [24] treat the non-parametric asymptotics of (even non-separable) tapers coherently. Here we follow [28] in producing estimators from the forthcoming spectral eq. (46), which removes bias from boundary effects, can be tapered, maintains the statistical efficiency of a spatial likelihood, and remains computationally competitive. Should the parameters of the model become sample-size dependent, no existing theory is available yet.
@@ 4.2 Finite sampled data: heuristics
For sampled data, [58] defined the likelihood of observing the data ${\mathcal{H}}(\mathbf{x})$ under the spectral model (12)–(13) parameterized by $\boldsymbol{\theta}$. This ‘blurred likelihood’ $\bar{{\mathcal{L}}}(\boldsymbol{\theta})$ instead of ${\mathcal{L}}(\boldsymbol{\theta})$, is given in terms of the Fourier coefficients of the gridded and windowed data, $H(\mathbf{k})$ in eq. (25), and of the blurred isotropic spectral density, $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ of eq. (30), summed over all wavenumbers in the grid (21), by
@@
Compare eq. (46) with eq. (39): the only difference is the acknowledgment of the spectral blurring effect of sampled data. Eq. (46) is the quantity that we maximize, under positivity constraints, for the parameter vector $\boldsymbol{\theta}$, thereby defining the quasi-maximum-likelihood estimate $\hat{\theta}$ to be obtained by solving for the ‘score’ vector ${\mbox{$\bar{\gamma}$}}(\boldsymbol{\theta})$ of numerical first derivatives of the blurred likelihood, in the sense
${\mbox{$\bar{\gamma}$}}({\mbox{$\hat{\theta}$}})=\mathbf{0}.$ (47)
@@
Satisfying eq. (47) to find $\hat{\theta}$, for example by an iterative function minimization routine, requires repeated evaluation of the spectral density (12) on the grid (21), with the blurring implemented convolutionally (on a refined and subsequently reinterpolated grid) or else exactly, as in eq. (30). Without entering into more details for now, Fig. 5 shows the results of a suite of experiments conducted to illustrate the performance of our numerical method that recovers each of the three Matérn parameters $(\sigma^{2},\nu,\rho)$, as a function of grid size (measured in terms of the correlation length $\pi\rho$). As the right-hand side panels of Fig. 4 showed for the variance estimate, the maximum-likelihood estimates are very well-behaved, from about the point where the grid size reaches a linear dimension of about $\pi\rho$. The procedure almost surely yields low-variance and practically unbiased estimates from a grid size of about $2\pi\rho$ onward, as Fig. 5 shows empirically.
Figure 5: Behavior of the relative error of the Whittle maximum-likelihood estimators of the Matérn parameters $(\sigma^{2},\nu,\rho)$ for the two sets of true values listed in the titles, conducted on square lattices growing in size from $M=N=2$ to $M=N=128$ pixels of size $\Delta x=\Delta y=10~\mathrm{km}$, quoting $M\Delta x$ in multiples of $\pi\rho$. Gray bars cover the 5th through 95th percentiles of the estimates for each set of up to 40 simulations. Black filled circles are the means of the estimates, computed over the 80th percentile of the sets for fields whose linear dimension $M\Delta x<2\pi\rho$, but over the full set of up to 40 estimates beyond that size. With growing grid size, the estimates reveal themselves to be less biased with shrinking variance. The blue line tracks the absolute value of the relative error multiplied by $\sqrt{MN}$, offset by 0.5-0.5 for visual clarity.
@@
The mathematical form and geometry of the blurred likelihood function are what, fundamentally, controls the observed behaviors. [58] only considered convolutional approaches to blurring, and were limited in their ability to acknowledge the spectral interaction induced by the applied data windows on the parameter covariance estimates. The first and second derivatives of the blurred likelihood are not simply obtained by substituting $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(k)$ for ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ in eqs (40) and (42), since the factors $m_{\theta}(k)$ that appear in the expressions for the score and the Hessian are analytical derivatives that involve the unblurred spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$. Replacing ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ by $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(k)$ in eq. (40) yields reasonable approximations for the likelihood derivatives, which compare relatively favorably to their numerical counterparts—especially for large sample sizes. Most numerical optimization routines will be able to maximize eq. (46), solving eq. (47) without being given explicit expressions for the score and the Hessian. However, in order to derive accurate confidence intervals on our estimated parameters, we do need access to the expected value of the second-order derivatives of the actual likelihood that is being maximized.
@@
Numerical experiments and theoretical considerations along the lines suggested in their Appendix A8 tempted [58] into concluding that eq. (45) could be used to construct confidence intervals for the solutions of eq. (47) in our present case of univariate two-dimensional analysis. Under the viewpoint espoused in their eqs (A84)–(A87), the blurred spectrum is an additive correction term (small for smoothly varying spectra) away from the original. In this framework, neglecting to blur the Fisher matrix (not to mention neglecting wavenumber correlations) was believed to have an altogether negligible effect on the estimation variances based on its inverse, even if blurring the likelihood is absolutely essential to arrive at the estimate in the first place. However, the ability of the unblurred Fisher matrix to help predict the variance of the parameters derived via maximization of the blurred likelihoods turns out to be poor, especially as concerns the variance and correlation parameters $\sigma^{2}$} and $\rho$. The unblurred expression, eq. (43), of the Fisher matrix provides an asymptotic but ultimately inadequate match to the average of the numerical Hessian for real-world sampling scenarios.
@@
Not accounting for wavenumber correlation proved to be another stumbling block. [58] conceived of approximations to account for wavenumber correlation involving a multiplicative correction term (their eqs A56–A58). For very large sample sizes this correction term approaches unity. Contrary to the optimism they expressed, uncertainty estimates for the maximizers of eq. (46) that rely on eq. (45) are inadequate for all but the largest sample sizes. A heuristic way of determining the estimation variance for the recovered parameters when actual data are being investigated is to generate synthetics with features identical to those of the gridded data, from models with Matérn parameters given by previously obtained solutions, then estimating their parameters a number of times, and learning from their distribution what the likely uncertainty ranges for the parameters of the actual data patches might be, as in Fig. 5. However justified, little transferable knowledge is gained in the process, and the procedure is cumbersome and time-consuming.
@@
[28] showed the way forward by further developing the theory of likelihood analysis for finite sampled data, on which we rely to develop the practical methods offered in the next sections. They include the ability to calculate uncertainty estimates on the parameters from first principles. The next section provides a complete description of the entire workflow.